OpenStudy (anonymous):

Find the difference quotient of f f(x+h)-f(x)/h, h does not equal 0 f(x)=-8x+4

OpenStudy (anonymous):

f(x)=-8x+4 f(x +h) = -8(x+h) +4 = -8x -8h +4 f(x+h)-f(x)/h = [ -8x -8h +4 -(-8x +4) ] / h = (-8x -8h +4 +8x -4) / h = -8h/h = -8

OpenStudy (anonymous):

\[f(x+h)=-8(x+h)+4=-8x-8h+4\] \[f(x+h)-f(x)=/8x-8h+4-(8x+4)=-8h\] \[\frac{f(x+h)-f(x)}{h}=\frac{8h}{h}=-8\] aka the slope

OpenStudy (anonymous):

lol those are your favorite problems satellite :D

OpenStudy (anonymous):

yeah i an noticed not one but two typos

OpenStudy (anonymous):

my all time favorite one of these to work out was \[f(x)=\sqrt[3]{x^2}\]